OpenStudy (uri):

Stats help Q.From five boys & 4 girls a committe of 5 is to be formed.In how many ways can this be done to include atleast one girl.

terenzreignz (terenzreignz):

I hope you remember what \[\Large ^nC_r\] means...

OpenStudy (uri):

It wasn't that...:D you thought me with different formula.

OpenStudy (uri):

But i know it's like nCr =nPr/n

terenzreignz (terenzreignz):

It's taught. And yes, I suppose so. Let's highlight the difference between \[\Large ^nP_r\] and\[\Large ^nC_r\]

OpenStudy (uri):

put numbers instead of N and R so i could actually tell you how..:D

terenzreignz (terenzreignz):

\[\Large ^nP_r\] counts how many ways you can arrange r elements in a row, from a set of n. On the other hand \[\Large ^nC_r\] counts how many ways you can CHOOSE r elements from a set of n. In other words, for \(\Large ^nP_r\), order matters, while in \(\Large ^nC_r\), order does not matter.

terenzreignz (terenzreignz):

Now in forming a committee, you don't exactly arrange the members in a row, and it isn't especially important how you arrange them, so you can bet that in questions such as "how many ways can a committee blah blah blah blah" that the key thing to remember is \(\Large ^nC_r\)

OpenStudy (uri):

All right xD

terenzreignz (terenzreignz):

So, how do you plan on going about this question? The additional complication is the fact that there must be at least one girl on the committee...

OpenStudy (uri):

4C1 and 5C1?

terenzreignz (terenzreignz):

4C1 = 4 5C1 = 5 That is trivial, and those don't really help you :) It turns out, it is incredibly difficult to directly count the number of ways you can form a committee of five, while making sure to include at least one girl :)

terenzreignz (terenzreignz):

Another way involves just counting the number of ways you can form a committee of 5 from these 9 choices... and that is just...?

terenzreignz (terenzreignz):

How many ways can you form a committee of 5 from a set of 9 members?

terenzreignz (terenzreignz):

(9 of course, since there are four girls and five boys)

OpenStudy (uri):

Hm so 9C5?

terenzreignz (terenzreignz):

That's right, so evaluate that, you get?

terenzreignz (terenzreignz):

NOTE: That is not the answer yet...

OpenStudy (anonymous):

It's not 9C5 - 1?

terenzreignz (terenzreignz):

@KevinOrr I was getting there :P

OpenStudy (anonymous):

Oh sorrz

terenzreignz (terenzreignz):

But that's 9P5 :P

OpenStudy (uri):

9C5 =9*8*7*6*5/5 right?

terenzreignz (terenzreignz):

Nope... \[\Large ^nC_r =\frac{\Large ^nP_r}{\color{red}{r!}} \]

OpenStudy (uri):

So,9C5=9P5/5! =9*8*7*6*5/5*4*3*2*1

terenzreignz (terenzreignz):

Yes... just simplify that nasty looking expression... first... before I tell you the rest :D

OpenStudy (uri):

15,120/120=126.

terenzreignz (terenzreignz):

slow down. We get 9C5 What do you want 9C4 for?

terenzreignz (terenzreignz):

And of course they'd be equal :/ \[\Large ^9C_5 = \frac{9!}{(9-5)!5!}=\frac{9!}{4!5!}=\frac{9!}{(9-4!)4!}=^9C_4\]

OpenStudy (uri):

Now what after that? :D

terenzreignz (terenzreignz):

Anyway, we have 126 ways to form a committee of 5, from these 9 members, right?

OpenStudy (uri):

Yess.

terenzreignz (terenzreignz):

But these also include those ways that don't include a single girl. In other words, ways that the question doesn't 'want' lol

OpenStudy (uri):

Okay i get it.. :3 then?

terenzreignz (terenzreignz):

Then... we take a way the number of ways you can form a committee of 5 WITHOUT including a single girl. How do we do that? How many ways can we form a committee of 5 without including girls?

OpenStudy (uri):

We have to include at least one girl the questions says,No?

terenzreignz (terenzreignz):

Yes. So from the total number of ways to form a committee of 5 out of the 9 members, we HAVE TO take away the number of ways to form a committee of 5 WITHOUT GIRLS.

terenzreignz (terenzreignz):

So how do you count the number of ways you can form a committee of 5 without girls?

OpenStudy (uri):

9P5?

terenzreignz (terenzreignz):

That's what we just did... 9P5 counts the number of ways you can form a committee if you choose from all 9 of them, but remember, no girls? How many of them are not girls?

OpenStudy (uri):

5.

terenzreignz (terenzreignz):

So, to form a committee of 5 without girls, we may only choose from those 5 boys... So that means?

OpenStudy (uri):

5P4?

terenzreignz (terenzreignz):

4? Why not 5? You're forming a committee of 5, not 4 -.-

OpenStudy (uri):

Yeah that's what I'm thinking won't it be 5P5?

terenzreignz (terenzreignz):

wait, no, 5C5

terenzreignz (terenzreignz):

Why did you use P all of a sudden? -.-

OpenStudy (uri):

How do i know when to use P or C? :o

terenzreignz (terenzreignz):

Wasn't that kind of the first thing I said? -.-

OpenStudy (uri):

Anyway,5C5=5*4*3*2*1/5*4*3*2*1 =120/120 =1.

terenzreignz (terenzreignz):

You use P when order matters. You uce C when it doesn't. And in forming committees, order doesn't matter.

terenzreignz (terenzreignz):

Okay, so 5C5 is 1 Means there is only one way to form a committee of 5 without any girls.

terenzreignz (terenzreignz):

You take that away from the TOTAL number of ways you can form a committee of 5 from the 9,and that's your answer.

OpenStudy (uri):

Sorry but i have to say this *how do we know when order matters or not?* sorry...Sir.

terenzreignz (terenzreignz):

Use your logic. But look for keywords... if you see the word 'arrange' or anything similar, then order matters.

terenzreignz (terenzreignz):

on the other hand, you see the word 'form' or 'choose' (though this is ambiguous, look out for 'arrange') then order probably doesn't matter

OpenStudy (uri):

Ah i see.. :3 SO the answer is 125 ways.

terenzreignz (terenzreignz):

Yes.

OpenStudy (uri):

Am i not your favourite student? :3

terenzreignz (terenzreignz):

That remains to be seen :|

OpenStudy (uri):

Seen,ah?

terenzreignz (terenzreignz):

Vamos a ver.

OpenStudy (uri):

Bien entonces ... Pero cuando?

terenzreignz (terenzreignz):

Si no sabes ahora, nunca sabrás.

OpenStudy (uri):

Eso es tocar .. pero no lo sé: P

terenzreignz (terenzreignz):

Pero tu sabes que yo tengo muchos estudiantes que he ayudado :)

terenzreignz (terenzreignz):

Tu eres unica una :3

OpenStudy (uri):

Das besondere? : D

terenzreignz (terenzreignz):

Wir werden sehen. Was denkst du?

OpenStudy (uri):

Ich denke, was Sie denken, und Sie wissen, was Sie denken, nein? : D

terenzreignz (terenzreignz):

Nein, ich glaube dass du kannst nicht immer denken was ich denke.

OpenStudy (uri):

Autsch ...

terenzreignz (terenzreignz):

Wenn du glaubst dass wir sind Freunde, sag doch 'du' und nicht 'Sie' Wir brauchen nicht sehr formal sein :)

OpenStudy (uri):

Acabo de copiar lo que dice la traducción :)